OpenStudy (dezandalf42714):

Please help I am confused on what to do here. Mariahs sunflower grew 18 2/3 inches in one week. Her tulip plant grew 3 3/4 in one week. How many more inches did the sunflower grow than the tulip?

OpenStudy (anonymous):

Change it to 56/3 - 15/4

OpenStudy (dezandalf42714):

OHHHH thanks so much!!

OpenStudy (anonymous):

Now find a common denominator and subtract

OpenStudy (dezandalf42714):

Ok

OpenStudy (dezandalf42714):

So you would do the LCD which is 12. then you would do 3 x 4 which is 12 and 56 x 4 which is 224/12. then 4 x 3 is 12 then 15 x 3 is 45/12 then subtract which that would 179/12. and 12 goes into 179 14 11/12. So the sunflower grew 14 11/12 more than the tulip. = 56/3 − 15/4 = ((56 × 4) − (15 × 3)) / (3 × 4) = (224 - 45) / 12 = 179/12 = 179/12 = 14 11/12

OpenStudy (anonymous):

That's correct!